# iclr2025_conference.bib

### v3

@@ 6 Conclusion
In this paper, we introduced a pointwise memorization metric to quantify memorization effects in DPMs. We provided a theoretical analysis of conditional memorization, offering a generalized definition of informative labels and clarifying that random labels can also be informative. We distinguish between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional diffusion models by constructing a surrogate condition. The key to this approach is training a time-dependent classifier using our TDKD technique. We empirically validate SIDE on subsets of the CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS. We aim for our work to enhance the understanding of memorization mechanisms in diffusion models and inspire further methods to mitigate memorization.
@@
Our work proposes a training data extraction method targeted at diffusion models and the effectiveness of our proposed method has far exceeded random methods. This may be exploited by potential malicious users to attack existing generative diffusion models. By effectively avoiding the introduction of additional or unnecessary informative label information into the training process, the risk of a diffusion model being attacked by data extraction can be greatly reduced. Therefore, the contribution of our work to the community outweighs the risks it poses, especially considering that the scale of our extraction method has not reached the same scale of state-of-the-art diffusion models. Moreover, the theoretical and experimental understandings established in this work can help the community build a deeper understanding of the memorization mechanism of diffusion models, and inspire better defense methods.

### v5

@@ 7 Conclusion
In this paper, we distinguish the difference between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional DPMs by constructing surrogate conditions. The core of SIDE lies in training a time-dependent classifier using TDKD technique. We empirically validate effectiveness of SIDE on subsets of CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS. In addition, we introduce a point-wise memorization metric to quantify memorization effects in DPMs for theoretical analysis. Finally, we introduce theoretical understanding of memorization in conditional generative models, offering a generalized definition of informative labels, clarifying random labels can also be informative and explaining why SIDE works. Our work aims to deepen understanding of memorization in DPMs and inspire further methods to mitigate memorization.
Supplementary Material
@@
Our work introduces a memorization metric that not only quantifies memorization effects in diffusion models but also extends to deep learning models more broadly. This contribution is significant in enhancing our understanding of when and how models memorize training data, which is critical for addressing concerns about data privacy and model robustness. By providing a theoretical framework for conditional memorization, we pave the way for developing effective memorization mitigation strategies tailored for diffusion models. These advancements can lead to the design of more secure and trustworthy AI systems, reducing the risk of potential data leakage while fostering greater accountability in the use of generative technologies. Ultimately, our findings aim to empower researchers and practitioners to create models that better respect privacy, thus benefiting the wider AI community.
Appendix A Proof of Theorem1
By the assumption $TV(p_{i},p_{\theta_{i}})=0$ in the theorem, we can replace $p_{\theta_{c}}$ with $p_{c}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness. Firstly, we compute $\mathcal{M}_{point}(\mathcal{D}_{c};p_{c})$ here: